# Anabelian aspects of the outer automorphism groups of the absolute Galois groups of mixed-characteristic local fields

### v1

@@ Introduction
Let $p$ be a prime number, $k$ a finite extension of $\mathbb{Q}_{p}$, and $\overline{k}$ an algebraic closure of $k$. We write $G_{k}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k}/k)$ for the absolute Galois group of $k$ determined by the algebraic closure $\overline{k}$ and $\operatorname{Out}(G_{k})$ for the group of outer automorphisms of the group $G_{k}$ [or, equivalently, the group of outer continuous automorphisms of the profinite group $G_{k}$ — cf. [Hsh3], Proposition 3.3, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(\hskip 0.80002ptii\hskip 0.80002pt)$]. In the present paper, we study the outer automorphism group $\operatorname{Out}(G_{k})$ from the point of view of anabelian geometry.
A natural and fundamental question is to what extent various kinds of data concerning a field $k$ can be recovered from the group $G_{k}$. In particular, traditional approaches in anabelian geometry have focused on the question of whether the functor “taking absolute Galois groups” is fully faithful. Unfortunately, this functor from the groupoid of fields isomorphic to finite extensions of $\mathbb{Q}_{p}$ equipped with algebraic closures to the groupoid of profinite groups is faithful but not full if $p$ is odd [cf., e.g., the discussion given at the final portion of [N-S-W], chapter Chapter VII, §5]. For example, if $p$ is an odd prime, and $k/\mathbb{Q}_{p}$ is Galois but not abelian, then there exists a pair $(k^{\prime},\overline{k^{\prime}})$, where $k^{\prime}$ is a finite extension of $\mathbb{Q}_{p}$, and $\overline{k^{\prime}}$ is the algebraic closure of $k^{\prime}$, such that $k$ and $k^{\prime}$ are not isomorphic as fields, but the absolute Galois groups $G_{k}$ and $G_{k^{\prime}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k^{\prime}}/k^{\prime})$ are isomorphic as [topological] groups [cf. [Hsh2], Theorem F, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$].
Write $\operatorname{Aut}(k)$ for the group of automorphisms of the field $k$. Then we have a natural injective homomorphism $\operatorname{Aut}(k)\hookrightarrow\operatorname{Out}(G_{k})$ of groups [cf., e.g., [Hsh3], Proposition 2.1]. In the present paper, let us regard $\operatorname{Aut}(k)$ as a [necessarily finite] subgroup of $\operatorname{Out}(G_{k})$ by means of this injective homomorphism:
@@
We refer to the subgroup as a field-theoretic subgroup. Here, we note that it is well-known [cf., e.g., the discussion given at the final portion of [N-S-W], Chapter VII, §5] that [although a similar equality always holds for a finite extension of $\mathbb{Q}$ by the Neukirch-Uchida theorem — cf., e.g., [N-S-W], Chapter XII, Corollary 12.2.2], in general, the equality $\operatorname{Aut}(k)=\operatorname{Out}(G_{k})$ does not hold. Thus, we are interested in the difference between $\operatorname{Aut}(k)$ and $\operatorname{Out}(G_{k})$.
Let us recall that various “ring-theoretic characterizations” of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ have been studied in works such as [Mzk1], [Mzk2], and [Hsh1].
Theorem A ([Mzk1], [Mzk2], [Hsh1]).
@@
As a certain refinement of some observations concerning the implication $(2)\Rightarrow(1)$ of Theorem A, the following theorem was proved in [Hsh4].
Theorem B ([Hsh4]).
@@
To prove Theorem B, it is important to show that there exists an automorphism $\varphi$ of $G_{k}$ such that the induced automorphism $\varphi^{\times}$ of $k^{\times}$, which is obtained via the mono-anabelian reconstruction algorithm, does not preserve the subgroup $\mathbb{Z}_{p}^{\times}\subset k^{\times}$ of $k^{\times}$. When $k/\mathbb{Q}_{p}$ is a quadratic extension, we construct such an automorphism of $G_{k}$ explicitly by applying the techniques developed in [H-N] [cf. Theorem D below].
@@
is a characteristic subgroup [cf. [H-N], Lemma 2.6, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$]. Thus, we obtain an automorphism $\varphi^{\prime}$ of $G_{k^{\prime}}$ such that $(\varphi^{\prime})^{\times}$ does not preserve the subgroup $\mathbb{Z}_{p}^{\times}\subset(k^{\prime})^{\times}$ of $(k^{\prime})^{\times}$. Moreover, the restriction $\varphi$ to $G_{k}$ of the automorphism $\varphi^{\prime}$ satisfies the desired property.
@@
We write $\widehat{k^{\times}}$ for the profinite completion of the multiplicative group of $k$ and $k_{+}$ for the $\mathbb{Q}_{p}$-vector space obtained by forming the underlying additive module of $k$. When $p$ is odd, we fix topological generators $\sigma$, $\tau$, $x_{0},\ldots,x_{d}$ of $G_{k}$ as in Theorem D. Let $F\colon G_{k}\to k_{+}$ be the homomorphism obtained as the composite of the natural surjection $G_{k}\twoheadrightarrow G_{k}^{\operatorname{ab}}$, the inverse of the isomorphism induced by the local reciprocity map $\widehat{k^{\times}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k}^{\operatorname{ab}}$, the projection $\widehat{k^{\times}}\twoheadrightarrow\mathcal{O}_{k}^{\times}$ determined by a choice of a uniformizer of $\mathcal{O}_{k}$, and the $p$-adic logarithm $\log_{k}\colon\mathcal{O}_{k}^{\times}\to k_{+}$. Moreover, if $p$ is odd, then, for each $i\in\{1,2,\ldots,d\}$, we write $y_{i}\in k_{+}$ for the image of $x_{i}$ by the map $F$. Note that, when $d>1$ and $p$ is odd, the $d$ elements $y_{1},\ldots,y_{d}$ form a basis of the $\mathbb{Q}_{p}$-vector space $k_{+}$ [cf. [H-N], Lemma 1.2; [H-N], Lemma 1.3; [Hsh3], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$].
We obtain the following result by considering the mono-anabelian reconstruction algorithm, together with the “Dehn twists” in $\operatorname{Out}(G_{k})$ [cf., e.g., [Kum], [Grp]]. This result is interesting in its own right and, moreover, plays an important role in the proof of Theorem C.
@@
We obtain a useful description of the sub-$\mathbb{Q}_{p}$-vector space $(\mathbb{Q}_{p})_{+}\subset k_{+}$ in terms of the basis $y_{1},\ldots,y_{d}$ by applying Theorem E. This description enables us to apply the techniques developed in [Hsh4] to prove Theorem C, which is the first main result of the present paper.
Here, we note that the generators and relations of $G_{k}$ in Theorem D are non-canonical, and that the proof of Theorem D is purely group-theoretic. Moreover, it is non-trivial to relate these generators and relations of $G_{k}$ to the ring structure of $k$. From this perspective, it is of independent interest the issue of whether or not, under rather mild assumptions such as those in Theorem E, one can obtain a result that directly connects the generators and relations of $G_{k}$ in Theorem D with the ring structure of $k$.
Next, let us discuss the second main theorem of the present paper. First, we recall that one interesting problem from the perspective of algorithmic anabelian geometry is the issue of whether or not the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ can be recovered functorially and group-theoretically from $G_{k}$. For example, as shown in [Hsh2], Theorem 6.12, under suitable assumptions on $k$, the set of $\operatorname{Out}(G_{k})$-conjugates of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ can be recovered functorially and group-theoretically.
Theorem F ([Hsh2]).
Suppose that $k$ is Galois-specifiable [cf. [Hsh2], Definition 6.1]. Then the set of $\operatorname{Out}(G_{k})$-conjugates of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is equal to the set of strictly quasi-geometric subgroups of $\operatorname{Out}(G_{k})$ [cf. [Hsh2], Definition 6.5, (i), (ii)]. In particular, the set of $\operatorname{Out}(G_{k})$-conjugates of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ can be recovered functorially and group-theoretically.
It follows formally that if one can establish a functorial group-theoretic algorithm that recovers the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ from $G_{k}$, then it is necessary that the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is a normal subgroup [cf. Remark 2.20]. However, the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is not necessarily a normal subgroup in general as shown in [Hsh2] and [H-N].
Theorem G ([Hsh2]).
@@
We apply essentially different arguments to prove Theorem I depending on whether $k\colon\mathbb{Q}_{p}]$ is even or odd. Suppose that $k\colon\mathbb{Q}_{p}]$ is even. Let $\varphi$ be an automorphism of $G_{k}$ as in Theorem C. We write $\varphi_{+}$ for the automorphisms, which is induced from $\varphi$ by the mono-anabelian reconstruction algorithm. Suppose that $k/\mathbb{Q}_{p}$ is a finite Galois extension. It follows from the discussion in §1 that the automorphism $\varphi_{+}$ does not preserve the $\mathbb{Q}_{p}$-subspace $(\mathbb{Q}_{p})_{+}\subset k_{+}$ of $k_{+}$. We prove Theorem I in the case where $k/\mathbb{Q}_{p}$ is a finite Galois extension of even degree by combining this result with the argument of [H-N].
On the other hand, when $k/\mathbb{Q}_{p}$ is of odd degree, since it is not clear to the author whether or not there exists an element of $\operatorname{Out}(G_{k})$ that does not preserve the $\mathbb{Q}_{p}$-subspace $(\mathbb{Q}_{p})_{+}\subset k_{+}$, the same technique cannot be applied. A key observation in the proof of Theorem I in the case where $k/\mathbb{Q}_{p}$ is a finite Galois extension of odd degree greater than one is obtained by focusing on the analogy between $\operatorname{Out}(G_{k})$ and the mapping class group of a surface, as developed in [Kum], [Grp], and related works.
In the present paper, we prove Theorem I by comparing, from the perspective of the natural $p$-adic Lie group structures, the image of the “subgroup of $\operatorname{Out}(G_{k})$ corresponding to the mapping class group” [cf. Remark 2.16] by the group homomorphism $\operatorname{Out}(G_{k})\to\operatorname{Aut}_{\mathbb{Q}_{p}}(k_{+})$ induced by the mono-anabelian reconstruction algorithm of [Hsh3], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$, with the image of $Z_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$ by the same homomorphism.
@@
Although [Kum] and [Grp] studied certain analogies between the outer automorphism groups of absolute Galois groups of mixed-characteristic local fields and the mapping class groups of topological surfaces, there have been very few results that apply this analogy to objects reflecting the ring structure of $k$. In the present paper, some results from the theory of mapping class groups are applied to prove the non-normality of field-theoretic subgroups and to establish nontrivial phenomena concerning $p$-adic representations. To the best of the author’s knowledge, this is the first work that applies results from the theory of mapping class groups to the anabelian geometry of mixed-characteristic local fields in a way that goes beyond a mere analogy.
0 Notational conventions
@@
Groups of MLF-type. We shall refer to a group isomorphic to the absolute Galois group of an MLF as a group of MLF-type. Let $G$ be a group of MLF-type. We shall say that a subgroup of $G$ is open if it is of finite index. It follows from [N-D], Theorem 0.1, and an elementary argument from general topology that the open subgroups of $G$ determine the structure of a profinite group of $G$, and that this topology of $G$ is a uniquely determined topology of $G$ that gives rise to a structure of profinite group on $G$. In the remainder of the present paper, we equip $G$ with this profinite group structure. In particular, an arbitrary isomorphism between groups of MLF-type as abstract groups is necessarily an isomorphism of profinite groups.
Moreover, let us recall [cf. [Hsh3], Definition 3.5, Proposition 3.6, Definition 3.10, Proposition 3.11] that there exist functorial group-theoretic algorithms for constructing, from a group of MLF-type $G$,
@@
Moreover, it follows from [Hsh3], Proposition 3.11, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$; [H-N], Lemma 1.2, that we have natural homomorphisms
@@
1 A group-theoretic description of the kernels of absolute trace maps and the application to the study of Aut-intrinsic Hodge-Tate representations
In the present §1, let $k$ be an MLF, $\overline{k}$ an algebraic closure of $k$, and $G$ a group of MLF-type. In the present §1, we show that the kernel of the trace map for the extension $k/k^{(d=1)}$ can be described “group-theoretically”. Moreover, by means of this group-theoretic description of the kernel of the trace map for $k/k^{(d=1)}$, together with classical techniques from $p$-adic representation theory, we show that if $p_{k}$ is odd, $d_{k}$ is even, and $k$ is an absolutely Galois MLF, i.e., if $k/k^{(d=1)}$ is a Galois extension, then there exist a $\mathbb{Q}_{p_{k}}$-vector space $V$ of dimension $d_{k}$, a continuous representation $\rho\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$ that is irreducible, abelian, and crystalline [hence also Hodge-Tate], and a group automorphism $\varphi\colon G_{k}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k}$ such that $\rho\circ\varphi$ is not Hodge-Tate. This result generalizes the main theorem of [Hsh4].
@@
This assertion follows from [N-S-W], Theorem 7.5.14, together with [Hsh3], Proposition 3.6. ∎
In the remainder of the present paper, we apply the notational conventions introduced in the statement of Theorem 1.1 in each of the situations in which $p(G)$ is assumed to be odd. Moreover, for each $i=1,2,\ldots,d(G)$, write $y_{i}\in k_{+}(G)$ for the image of $x_{i}$ in $k_{+}(G)$ by the composite $P(G)\hookrightarrow I(G)\to\mathcal{O}^{\times}(G)\to k_{+}(G)$ [cf. conditions (1), (2) of Theorem 1.1; [Hsh3], Definition 3.10, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(\hskip 0.80002ptii\hskip 0.80002pt)$, $\rm(\hskip 0.59998ptv\hskip 0.59998pt)$].
We recall that the topological group $k_{+}(G)$ has a natural structure of $\mathbb{Q}_{p(G)}$-vector space of dimension $d(G)$ [cf. [H-N], Lemma 1.2]. Moreover, one verifies easily that the isomorphism of topological groups $k_{+}(G_{k})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}k_{+}$ of [Hsh3], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$ is also an isomorphism of $\mathbb{Q}_{p_{k}}$-vector spaces [here, we have $p_{k}=p(G_{k})$ — cf. [Hsh3], Proposition 3.6]. When $d(G)>1$ and $p(G)$ is odd, the $d(G)$ elements $y_{1},\ldots,y_{d(G)}$ defined in the preceding paragraph form a basis of the $\mathbb{Q}_{p(G)}$-vector space $k_{+}(G)$ [cf. [H-N], Lemma 1.3].
@@
This assertion follows immediately from [H-N], Lemma 2.3, $\rm(\hskip 0.80002ptii\hskip 0.80002pt)$. ∎
@@
where $\langle S\rangle_{\mathbb{Q}_{p_{k}}}$” denotes the sub-$\mathbb{Q}_{p_{k}}$-vector space of $k_{+}$ generated by “SS”. Here, by abuse of notation, for each integer $i$ satisfying $1\leq i\leq d_{k}$, we write $y_{i}\in k_{+}$ for the image of $y_{i}\in k_{+}(G_{k})$ by the isomorphism $k_{+}(G_{k})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}k_{+}$ of [Hsh3], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$.

### v2

@@ Abstract.
Key words and phrases:
anabelian geometry, mixed-characteristic local field, absolute Galois group, outer automorphism group, $p$-adic Hodge theory, mapping class group
2020 Mathematics Subject Classification:
11S20, 11S31, 11F80
@@ Introduction
Let $p$ be a prime number, $k$ a finite extension of $\mathbb{Q}_{p}$, and $\overline{k}$ an algebraic closure of $k$. We write $G_{k}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k}/k)$ for the absolute Galois group of $k$ determined by the algebraic closure $\overline{k}$ and $\operatorname{Out}(G_{k})$ for the group of outer automorphisms of the group $G_{k}$ [or, equivalently, the group of outer continuous automorphisms of the profinite group $G_{k}$ — cf. [8], Proposition 3.3, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(\hskip 0.80002ptii\hskip 0.80002pt)$]. In the present paper, we study the outer automorphism group $\operatorname{Out}(G_{k})$ from the point of view of anabelian geometry.
A natural and fundamental question is to what extent various kinds of data concerning a field $k$ can be recovered from the group $G_{k}$. In particular, traditional approaches in anabelian geometry have focused on the question of whether the functor “taking absolute Galois groups” is fully faithful. Unfortunately, this functor from the groupoid of fields isomorphic to finite extensions of $\mathbb{Q}_{p}$ equipped with algebraic closures to the groupoid of profinite groups is faithful but not full if $p$ is odd [cf., e.g., the discussion given at the final portion of [15], chapter Chapter VII, §5]. For example, if $p$ is an odd prime, and $k/\mathbb{Q}_{p}$ is Galois but not abelian, then there exists a pair $(k^{\prime},\overline{k^{\prime}})$, where $k^{\prime}$ is a finite extension of $\mathbb{Q}_{p}$, and $\overline{k^{\prime}}$ is the algebraic closure of $k^{\prime}$, such that $k$ and $k^{\prime}$ are not isomorphic as fields, but the absolute Galois groups $G_{k}$ and $G_{k^{\prime}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k^{\prime}}/k^{\prime})$ are isomorphic as [topological] groups [cf. [7], Theorem F, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$].
Write $\operatorname{Aut}(k)$ for the group of automorphisms of the field $k$. Then we have a natural injective homomorphism $\operatorname{Aut}(k)\hookrightarrow\operatorname{Out}(G_{k})$ of groups [cf., e.g., [8], Proposition 2.1]. In the present paper, let us regard $\operatorname{Aut}(k)$ as a [necessarily finite] subgroup of $\operatorname{Out}(G_{k})$ by means of this injective homomorphism:
@@
We refer to the subgroup as a field-theoretic subgroup. Here, we note that it is well-known [cf., e.g., the discussion given at the final portion of [15], Chapter VII, §5] that [although a similar equality always holds for a finite extension of $\mathbb{Q}$ by the Neukirch-Uchida theorem — cf., e.g., [15], Chapter XII, Corollary 12.2.2], in general, the equality $\operatorname{Aut}(k)=\operatorname{Out}(G_{k})$ does not hold. Thus, we are interested in the difference between $\operatorname{Aut}(k)$ and $\operatorname{Out}(G_{k})$.
Let us recall that various “ring-theoretic characterizations” of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ have been studied in works such as [13], [14], and [6].
Theorem A ([13], [14], [6]).
@@
As a certain refinement of some observations concerning the implication $(2)\Rightarrow(1)$ of Theorem A, the following theorem was proved in [9].
Theorem B ([9]).
@@
To prove Theorem B, it is important to show that there exists an automorphism $\varphi$ of $G_{k}$ such that the induced automorphism $\varphi^{\times}$ of $k^{\times}$, which is obtained via the mono-anabelian reconstruction algorithm, does not preserve the subgroup $\mathbb{Z}_{p}^{\times}\subset k^{\times}$ of $k^{\times}$. When $k/\mathbb{Q}_{p}$ is a quadratic extension, we construct such an automorphism of $G_{k}$ explicitly by applying the techniques developed in [10] [cf. Theorem D below].
@@
is a characteristic subgroup [cf. [10], Lemma 2.6, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$]. Thus, we obtain an automorphism $\varphi^{\prime}$ of $G_{k^{\prime}}$ such that $(\varphi^{\prime})^{\times}$ does not preserve the subgroup $\mathbb{Z}_{p}^{\times}\subset(k^{\prime})^{\times}$ of $(k^{\prime})^{\times}$. Moreover, the restriction $\varphi$ to $G_{k}$ of the automorphism $\varphi^{\prime}$ satisfies the desired property.
@@
We write $\widehat{k^{\times}}$ for the profinite completion of the multiplicative group of $k$ and $k_{+}$ for the $\mathbb{Q}_{p}$-vector space obtained by forming the underlying additive module of $k$. When $p$ is odd, we fix topological generators $\sigma$, $\tau$, $x_{0},\ldots,x_{d}$ of $G_{k}$ as in Theorem D. Let $F\colon G_{k}\to k_{+}$ be the homomorphism obtained as the composite of the natural surjection $G_{k}\twoheadrightarrow G_{k}^{\operatorname{ab}}$, the inverse of the isomorphism induced by the local reciprocity map $\widehat{k^{\times}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k}^{\operatorname{ab}}$, the projection $\widehat{k^{\times}}\twoheadrightarrow\mathcal{O}_{k}^{\times}$ determined by a choice of a uniformizer of $\mathcal{O}_{k}$, and the $p$-adic logarithm $\log_{k}\colon\mathcal{O}_{k}^{\times}\to k_{+}$. Moreover, if $p$ is odd, then, for each $i\in\{1,2,\ldots,d\}$, we write $y_{i}\in k_{+}$ for the image of $x_{i}$ by the map $F$. Note that, when $d>1$ and $p$ is odd, the $d$ elements $y_{1},\ldots,y_{d}$ form a basis of the $\mathbb{Q}_{p}$-vector space $k_{+}$ [cf. [10], Lemma 1.2; [10], Lemma 1.3; [8], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$].
We obtain the following result by considering the mono-anabelian reconstruction algorithm, together with the “Dehn twists” in $\operatorname{Out}(G_{k})$ [cf., e.g., [12], [4]]. This result is interesting in its own right and, moreover, plays an important role in the proof of Theorem C.
@@
We obtain a useful description of the sub-$\mathbb{Q}_{p}$-vector space $(\mathbb{Q}_{p})_{+}\subset k_{+}$ in terms of the basis $y_{1},\ldots,y_{d}$ by applying Theorem E. This description enables us to apply the techniques developed in [9] to prove Theorem C, which is the first main result of the present paper.
Here, we note that the generators and relations of $G_{k}$ in Theorem D are non-canonical, and that the proof of Theorem D is purely group-theoretic. Moreover, it is nontrivial to relate these generators and relations of $G_{k}$ to the ring structure of $k$. From this perspective, it is of independent interest the issue of whether or not, under rather mild assumptions such as those in Theorem E, one can obtain a result that directly connects the generators and relations of $G_{k}$ in Theorem D with the ring structure of $k$.
Next, let us discuss the second main theorem of the present paper. First, we recall that one interesting problem from the perspective of algorithmic anabelian geometry is the issue of whether or not the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ can be recovered functorially and group-theoretically from $G_{k}$. For example, as shown in [7], Theorem 6.12, under suitable assumptions on $k$, the set of $\operatorname{Out}(G_{k})$-conjugates of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ can be recovered functorially and group-theoretically.
Theorem F ([7]).
Suppose that $k$ is Galois-specifiable [cf. [7], Definition 6.1]. Then the set of $\operatorname{Out}(G_{k})$-conjugates of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is equal to the set of strictly quasi-geometric subgroups of $\operatorname{Out}(G_{k})$ [cf. [7], Definition 6.5, (i), (ii)]. In particular, the set of $\operatorname{Out}(G_{k})$-conjugates of the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ can be recovered functorially and group-theoretically.
It follows formally that if one can establish a functorial group-theoretic algorithm that recovers the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ from $G_{k}$, then it is necessary that the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is a normal subgroup [cf. Remark 2.20]. However, the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is not necessarily a normal subgroup in general as shown in [7] and [10].
Theorem G ([7]).
@@
We apply essentially different arguments to prove Theorem I depending on whether $k\colon\mathbb{Q}_{p}]$ is even or odd. Suppose that $k\colon\mathbb{Q}_{p}]$ is even. Let $\varphi$ be an automorphism of $G_{k}$ as in Theorem C. We write $\varphi_{+}$ for the automorphisms, which is induced from $\varphi$ by the mono-anabelian reconstruction algorithm. Suppose that $k/\mathbb{Q}_{p}$ is a finite Galois extension. It follows from the discussion in §1 that the automorphism $\varphi_{+}$ does not preserve the $\mathbb{Q}_{p}$-subspace $(\mathbb{Q}_{p})_{+}\subset k_{+}$ of $k_{+}$. We prove Theorem I in the case where $k/\mathbb{Q}_{p}$ is a finite Galois extension of even degree by combining this result with the argument of [10].
On the other hand, when $k/\mathbb{Q}_{p}$ is of odd degree, since it is not clear to the author whether or not there exists an element of $\operatorname{Out}(G_{k})$ that does not preserve the $\mathbb{Q}_{p}$-subspace $(\mathbb{Q}_{p})_{+}\subset k_{+}$, the same technique cannot be applied. A key observation in the proof of Theorem I in the case where $k/\mathbb{Q}_{p}$ is a finite Galois extension of odd degree greater than one is obtained by focusing on the analogy between $\operatorname{Out}(G_{k})$ and the mapping class group of a surface, as developed in [12], [4], and related works.
In the present paper, we prove Theorem I by comparing, from the perspective of the natural $p$-adic Lie group structures, the image of the “subgroup of $\operatorname{Out}(G_{k})$ corresponding to the mapping class group” [cf. Remark 2.16] by the group homomorphism $\operatorname{Out}(G_{k})\to\operatorname{Aut}_{\mathbb{Q}_{p}}(k_{+})$ induced by the mono-anabelian reconstruction algorithm of [8], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$, with the image of $Z_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$ by the same homomorphism.
@@
Although [12] and [4] studied certain analogies between the outer automorphism groups of absolute Galois groups of mixed-characteristic local fields and the mapping class groups of topological surfaces, there have been very few results that apply this analogy to objects reflecting the ring structure of $k$. In the present paper, some results from the theory of mapping class groups are applied to prove the non-normality of field-theoretic subgroups and to establish nontrivial phenomena concerning $p$-adic representations. To the best of the author’s knowledge, this is the first work that applies results from the theory of mapping class groups to the anabelian geometry of mixed-characteristic local fields in a way that goes beyond a mere analogy.
0. Notational conventions
@@
Groups of MLF-type. We shall refer to a group isomorphic to the absolute Galois group of an MLF as a group of MLF-type. Let $G$ be a group of MLF-type. We shall say that a subgroup of $G$ is open if it is of finite index. It follows from [16], Theorem 0.1, and an elementary argument from general topology that the open subgroups of $G$ determine the structure of a profinite group of $G$, and that this topology of $G$ is a uniquely determined topology of $G$ that gives rise to a structure of profinite group on $G$. In the remainder of the present paper, we equip $G$ with this profinite group structure. In particular, an arbitrary isomorphism between groups of MLF-type as abstract groups is necessarily an isomorphism of profinite groups.
Moreover, let us recall [cf. [8], Definition 3.5, Proposition 3.6, Definition 3.10, Proposition 3.11] that there exist functorial group-theoretic algorithms for constructing, from a group of MLF-type $G$,
@@
Moreover, it follows from [8], Proposition 3.11, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$; [10], Lemma 1.2, that we have natural homomorphisms
@@
1. A group-theoretic description of the kernels of absolute trace maps and the application to the study of Aut-intrinsic Hodge-Tate representations
In the present §1, let $k$ be an MLF, $\overline{k}$ an algebraic closure of $k$, and $G$ a group of MLF-type. In the present §1, we show that the kernel of the trace map for the extension $k/k^{(d=1)}$ can be described “group-theoretically”. Moreover, by means of this group-theoretic description of the kernel of the trace map for $k/k^{(d=1)}$, together with classical techniques from $p$-adic representation theory, we show that if $p_{k}$ is odd, $d_{k}$ is even, and $k$ is an absolutely Galois MLF, i.e., if $k/k^{(d=1)}$ is a Galois extension, then there exist a $\mathbb{Q}_{p_{k}}$-vector space $V$ of dimension $d_{k}$, a continuous representation $\rho\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$ that is irreducible, abelian, and crystalline [hence also Hodge-Tate], and a group automorphism $\varphi\colon G_{k}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k}$ such that $\rho\circ\varphi$ is not Hodge-Tate. This result generalizes the main theorem of [9].
@@
This assertion follows from [15], Theorem 7.5.14, together with [8], Proposition 3.6. ∎
In the remainder of the present paper, we apply the notational conventions introduced in the statement of Theorem 1.1 in each of the situations in which $p(G)$ is assumed to be odd. Moreover, for each $i=1,2,\ldots,d(G)$, write $y_{i}\in k_{+}(G)$ for the image of $x_{i}$ in $k_{+}(G)$ by the composite $P(G)\hookrightarrow I(G)\to\mathcal{O}^{\times}(G)\to k_{+}(G)$ [cf. conditions (1), (2) of Theorem 1.1; [8], Definition 3.10, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(\hskip 0.80002ptii\hskip 0.80002pt)$, $\rm(\hskip 0.59998ptv\hskip 0.59998pt)$].
We recall that the topological group $k_{+}(G)$ has a natural structure of $\mathbb{Q}_{p(G)}$-vector space of dimension $d(G)$ [cf. [10], Lemma 1.2]. Moreover, one verifies easily that the isomorphism of topological groups $k_{+}(G_{k})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}k_{+}$ of [8], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$ is also an isomorphism of $\mathbb{Q}_{p_{k}}$-vector spaces [here, we have $p_{k}=p(G_{k})$ — cf. [8], Proposition 3.6]. When $d(G)>1$ and $p(G)$ is odd, the $d(G)$ elements $y_{1},\ldots,y_{d(G)}$ defined in the preceding paragraph form a basis of the $\mathbb{Q}_{p(G)}$-vector space $k_{+}(G)$ [cf. [10], Lemma 1.3].
@@
This assertion follows immediately from [10], Lemma 2.3, $\rm(\hskip 0.80002ptii\hskip 0.80002pt)$. ∎
@@
where $\langle S\rangle_{\mathbb{Q}_{p_{k}}}$” denotes the sub-$\mathbb{Q}_{p_{k}}$-vector space of $k_{+}$ generated by “SS”. Here, by abuse of notation, for each integer $i$ satisfying $1\leq i\leq d_{k}$, we write $y_{i}\in k_{+}$ for the image of $y_{i}\in k_{+}(G_{k})$ by the isomorphism $k_{+}(G_{k})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}k_{+}$ of [8], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$.